# PARCO: Learning Parallel Autoregressive Policies for Efficient Multi-Agent Combinatorial Optimization

### v1

@@ B.1.3 PARCO Network Hyperparameters
Initial Embedding. This layer projects initial raw features to hidden space. For the depot, the initial embedding is the positional encoding of the depot’s location $X_{0}$. For agents, the initial embedding is the encoding for the initial location, capacity, and speed. Main Encoder. we employ $L=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP.
@@
Context Embedding. This layer projects dynamic raw features to hidden space. The context is the embedding for the depot states, current node states, current time, remaining capacities, time of backing to the depot, and number of visited nodes. These features are then employed to update multiple queries $q_{k},~{}k=1,\dots,m$ simultaneously. Main Decoder. Similarly to the encoder, we employ the same hidden dimension and number of attention heads for the Multiple Pointer Mechanism.
@@
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP. Unlike the encoder layer, which acts between all $m+N$ nodes, communication layers are lighter because they communicate between $m$ agents.
@@
We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action $a$ is the highest. In other words, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
@@ B.1.4 PARCO Training Hyperparameters
For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions. We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=10$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma and Ba, 2014) with a total batch size $512$ (using 4 GPUs in Distributed Data Parallel configuration) and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $4\times 10^{5}$ randomly generated data. Training takes around 15 hours in our configuration.
@@ B.2 OMDCPDP
B.2.1 Datasets
@@
Neural baselines were trained with the specific number of nodes $2N$ and number of agents $m$ they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $2N\sim\mathcal{U}(50,100)$ and $m\sim\mathcal{U}(10,50)$. As we show in the Table 2, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,2N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. Similarly, we sample $m$ initial vehicle locations from the same distribution. We set the demand $d_{i}$ for each customer to $1$ and the capacity of each vehicle to $3$. This emulates realistic settings in which a single package per customer will be picked up and delivered.
@@
Testing is performed on 100 new instances for each setting of $N$ and $m$ in Table 2 with the distributions from the training settings. We finally test PARCO on large-scale, real-world data from complex, realistic distributions with $32$ instances for each vehicle, pickup, and delivery location in Seoul City, South Korea. This dataset is provided in our source code. Fig. 4 provides an example of such data. In Table 2, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward).
Figure 4: Real-world instance for the OMDCPDP problem in Seoul City, South Korea, with $N=1000$ locations and $m=100$ agents (green) showing relations (gray) of pickups (red triangles) and their respective deliveries (black crosses).
B.2.2 Baselines
OR-Tools
The OR-Tools (Furnon and Perron, 2024) is open-source software designed to address various combinatorial optimization problems. This toolkit offers a comprehensive suite of solvers suitable for linear programming, mixed-integer programming, constraint programming, and routing and scheduling challenges. Specifically for routing problems like the OMDCPDP, OR-Tools can integrate additional constraints to enhance solution accuracy. Documentation for the Python OR-Tools library, along with example codes for diverse optimization challenges, is readily accessible on their official website. For our experiments, we maintained consistent parameters across various problem sizes and numbers of agents. We configured the global span cost coefficient to $10,000$, selected PATH_CHEAPEST_ARC as the initial solution strategy, followed by GUIDED_LOCAL_SEARCH for local optimization. The solving time was capped at 300 seconds to standardize comparison and measure efficiency.
HAM
The Heterogeneous Attention Model (HAM) (Li et al., 2021a) utilizes a neural network integrated with a heterogeneous attention mechanism that distinguishes between the roles of nodes and enforces precedence constraints, ensuring the correct sequence of pickup and delivery nodes. This approach helps the deep reinforcement learning model to make informed node selections during route planning. We used the same hyperparameter settings from the original HAM experiments.
MAPDP
The Multi-Agent Reinforcement Learning-based Framework for Cooperative Pickup and Delivery Problem (MAPDP) (Zong et al., 2022) pioneers the use of multi-agent reinforcement learning (MARL) for the cooperative PDP with multiple vehicle agents. This innovative framework introduces a centralized MARL architecture to generate cooperative decisions among agents, incorporating a paired context embedding to capture the inter-dependency of heterogeneous nodes. We adapted the MAPDP to fit our OMDCPDP task, utilizing the same encoder as PARCO to ensure a fair comparison. For the decoder and training phases, we kept the same random conflict handler and retained the same hyperparameters as detailed in the original study with 8 heads, 128 hidden dimensions for model structure; $\mathrm{lr}=10^{-3},L_{\mathrm{Adam}}=3$ for turning.
@@ B.2.3 PARCO Network Hyperparameters
Most hyperparameters are kept similar to Section B.1.3; we report all of them nonetheless below.
@@
Initial Embedding. This layer projects initial raw features to hidden space. For depots, the initial embeddings encode the location $o_{k}$ and the respective vehicle’s capacity $Q_{k}$. For pickup nodes, the initial embeddings encode the location and paired delivery nodes’ location. For delivery nodes, the initial embeddings encode the location and paired pickup nodes’ location. Main Encoder. We employ $l=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP.
@@
Context Embedding. This layer projects dynamic raw features to hidden space. The context is the embedding for the depot states $o_{k}$, current node states, current length, remaining capacity, and number of visited nodes. These features are then employed to update multiple queries $q_{k},~{}k=1,\dots,m$ simultaneously. Main Decoder. Similarly to the encoder, we employ the same hidden dimension and number of attention heads for the Multiple Pointer Mechanism.
@@
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP. Note that unlike the encoder layer, which acts between all $m+N$ nodes, communication layers are lighter because they communicate between $m$ agents.
@@
We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action $a$ is the highest. In other words, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
@@ B.2.4 PARCO Training Hyperparameters
For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions. We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=8$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma and Ba, 2014) with a total batch size $128$ on a single GPU and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $10^{5}$ randomly generated data. Training takes less than 4 hours in our configuration.
@@ B.3.1 Baselines
The comparison of PARCO with other FFSP baselines is shown in table Table 3. Being a neural solver, we benchmark PARCO mainly against MatNet (Kwon et al., 2021), the SotA NCO architecture for the FFSP. MatNet is an encoder-decoder architecture, which is inspired by the attention model (Kool et al., 2018). It extends the encoder of the attention model with a dual graph attention layer, a horizontal stack of two transformer blocks, capable to encode nodes of different types in a bipartite graph - like machines and jobs in the FFSP. Kwon et al. (2021) train MatNet using POMO (Kwon et al., 2020).
CPLEX
Besides the aforementioned MatNet model, we implement the mathematical model described above in the exact solver CPLEX with a time budget of 60 and 600 seconds per instance. However, with both time budgets, CPLEX is only capable of generating solutions to the FFSP20 instances.
Random and Shortest Job First
Third, we use different (meta-)heuristics to benchmark PARCO on the FFSP. The Random and Shortest Job First (SJF) heuristics are simple construction strategies that build valid schedules in an iterative manner. Starting from an empty schedule, the Random construction heuristic iterates through time steps $t=0,\dots T$ and stages $i=1\dots S$ and randomly assigns jobs available at the given time to an idle machine of the respective stage until all jobs are scheduled. Likewise, the SJF proceeds by assigning job-machine pairs with the shortest processing time first.
Genetic Algorithm
The Genetic Algorithms (GA) are metaheuristics widely used by the OR community to tackle the FFSP (Kahraman et al., 2008). The GA iteratively improves multiple candidate solutions called chromosomes. Each chromosome consists of $S\times N$ real numbers, where $S$ is the number of stages and $N$ is the number of jobs. For each job at each stage, the integer part of the corresponding number indicates the assigned machine index, while the fractional part determines job priority when multiple jobs are available simultaneously. Child chromosomes are created through crossover, inheriting integer and fractional parts independently from two parents. Mutations, applied with a 30% chance, use one of four randomly selected methods: exchange, inverse, insert, or change. The implementation uses 25 chromosomes. One initial chromosome is set to the Shortest Job First (SJF) heuristic solution and the best-performing chromosome is preserved across iterations. Each instance runs for 1,000 iterations.
Particle Swarm Optimization
Finally, Particle Swarm Optimization (PSO) iteratively updates multiple candidate solutions called particles, which are updated by the weighted sum of the inertial value, the local best, and the global best at each iteration (Singh and Mahapatra, 2012). In this implementation, particles use the same representation as GA chromosomes. The algorithm employs 25 particles, with an inertial weight of 0.7 and cognitive and social constants set to 1.5. One initial particle represents the SJF heuristic solution. Like GA, PSO runs for 1,000 iterations per instance.
@@ B.3.2 Datasets
We follow the instance generation scheme outlined in Kwon et al. (2021) sample processing times for job-machine pairs independently from a uniform distribution within the bounds $[2,10]$. For the main experiments on the FFSP shown in Table 3, we also use the same instance sizes as Kwon et al. (2021) with $N=20,50\text{ and }100$ jobs (denoted FFSP20, FFSP50 and FFSP100, respectively) that need to be processed in $S=3$ stages on $m_{i}=4$ machines each (i.e. $m=12$ machines in total). For the agent sensitivity analysis shown in Table 4, we fix the number of jobs to 50 but alter the number of agents. Still, we use $S=3$ for this experiment, but alter the number of machines per stage to $m_{i}=6,8\text{ and }10$, yielding a total of 18, 24 and 30 agents, respectively.
@@
Testing is performed on 1,000 test instances generated according to the above sampling scheme. We use the same instances as Kwon et al. (2021) for the experiments of Table 3 and generate new ones for the experiments of Table 4.
@@ B.3.3 PARCO Network Hyperparameters
The machines are regarded as the agents in our PARCO framework. As such, their embeddings are used as queries $Q_{t}^{s}$ in the Multiple Pointer Mechanism Eq. 5, while job embeddings are used as the keys and values. In each decoding step, the machine embeddings are fused with a projection of the time the respective machine becomes idle. Similarly, job embeddings are augmented with a linear transformation of the time they become available in the respective stage before entering the attention head in Eq. 4.
@@
We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action is the highest. Formally, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
@@ B.3.4 PARCO Training Hyperparameters
Regarding the training setup, each training instance $i$ is augmented by a factor of 24, and the average makespan over the augmented instances is used as a shared baseline $b^{\text{shared}}_{i}$ for the REINFORCE gradient estimator of Eq. 9. We use the ADAM optimizer (Kingma and Ba, 2014) with a learning rate of $4\times 10^{-4}$, which we alter during training using a cosine annealing scheme. We train separate models for FFSP20, FFSP50 and FFSP100 instances for 100, 150 and 200 epochs, respectively, using 1,000 randomly generated instances per epoch split into batches of size 50.111Note: to avoid OOMs, For FFSP100 instances, batches are further split into mini-batches of size 25 whose gradients are accumulated. For the learned methods, the times given in parenthesis indicate the time required to train the respective models, whereas the other time information indicates the time required to solve a single instance. For the agent sensitivity analysis in Table 4, Matnet, as well as our PARCO model, are trained for 20 epochs on 1,000 instances per epoch. Duration in this table indicates the time required for training the respective models on the hardware specified in Section B.4.
@@
The following figures visualize the decoding for the machines of a given stage using MatNet and PARCO. As one can see in Fig. 6, MatNet requires a decoder forward pass for each machine to schedule a job on each of them. In contrast, as detailed in Fig. 6, PARCO can schedule jobs on all machines simultaneously through its Multiple Pointer Mechanism and Agent Handler, leading to significant efficiency gains.
Figure 5: A FFSP Decoding Step with MatNet
Figure 6: A FFSP Decoding Step with PARCO

### v2

@@ B.1.3 PARCO Network Hyperparameters
Initial Embedding. This layer projects initial raw features to hidden space. For the depot, the initial embedding is the positional encoding of the depot’s location $X_{0}$. For agents, the initial embedding is the encoding for the initial location, capacity, and speed. Main Encoder. we employ $L=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
@@
Context Embedding. This layer projects dynamic raw features to hidden space. The context is the embedding for the depot states, current node states, current time, remaining capacities, time of backing to the depot, and number of visited nodes. Multiple Pointer Mechanism. Similarly to the encoder, we employ the same hidden dimension and number of attention heads for the Multiple Pointer Mechanism.
@@
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP. Unlike the encoder layer, which acts between all $M+N$ problem tokens, communication layers are lighter because they communicate between ${M}$ agents.
@@
We use the Priority-based Conflict Handler guided by the model output probability for managing conflicts: priority is given to the agent whose probability of selecting the conflicting action is the highest (see Section 4.6).
@@ B.1.4 PARCO Training Hyperparameters
Unlike baselines, which are trained and tested on the same distribution, we train a single PARCO model that can effectively generalize over multiple size and agent distributions thanks to our flexible structure. We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=10$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma & Ba, 2014) with a total batch size $512$ (using 4 GPUs in Distributed Data Parallel configuration) and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $4\times 10^{5}$ randomly generated data. Training takes around 15 hours in our configuration.
@@ B.2 OMDCPDP
The setting introduced in OMDCPDP is a more general and realistic setting than the one introduced in Zong et al. (2022), particularly due to the multiple depots and the global lateness objective function which is harder to optimize than vanilla min-sum.
B.2.1 Baselines
OR-Tools
Google OR-Tools (Furnon & Perron, 2024) is an open-source software suite designed to address various combinatorial optimization problems. This toolkit offers a comprehensive selection of solvers suitable for linear programming, mixed-integer programming, constraint programming, and routing and scheduling challenges. Specifically for routing problems like the OMDCPDP, OR-Tools can integrate additional constraints to enhance solution accuracy. For our experiments, we maintained consistent parameters across various problem sizes and numbers of agents. We configured the global span cost coefficient to $10,000$, selected PATH_CHEAPEST_ARC as the initial solution strategy, followed by GUIDED_LOCAL_SEARCH for local optimization. The solving time was set as $\{30,60,300,600\}$ seconds for $N=\{50,100,500,1000\}$, respectively.
HAM
The Heterogeneous Attention Model (HAM) (Li et al., 2021a) utilizes a neural network-integrated with a heterogeneous attention mechanism that distinguishes between the roles of nodes and enforces precedence constraints, ensuring the correct sequence of pickup and delivery nodes. This approach helps the deep reinforcement learning model to make informed node selections during route planning. We adapt the original model to handle OMDCPDP with multiple agents akin to the sequential planning of Son et al. (2024).
MAPDP
The Multi-Agent Reinforcement Learning-based Framework for Cooperative Pickup and Delivery Problem (MAPDP) (Zong et al., 2022) introduces a cooperative PDP with multiple vehicle agents. This framework is trained via a centralized (MA)RL architecture to generate cooperative decisions among agents, incorporating a paired context embedding to capture the inter-dependency of heterogeneous nodes. We adapted the MAPDP to fit our OMDCPDP task, utilizing the same encoder as PARCO to ensure a fair comparison. For the decoder and training phases, we kept the same random conflict handler, and we retained the hyperparameters detailed in the original study overall.
B.2.2 Datasets
@@
Neural baselines were trained with the specific number of nodes $N$ and number of agents ${M}$ they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(50,100)$ and $m\sim\mathcal{U}(5,20)$. As we show in the Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. Similarly, we sample ${M}$ initial vehicle locations from the same distribution. We set the demand $d_{i}$ for each customer to $1$ and the capacity of each vehicle to $3$. This emulates realistic settings in which a single package per customer will be picked up and delivered.
@@
Testing is performed on $1000$ new instances for each setting of in-distribution $N$ and ${M}$ in Table 1 with the distributions from the training settings. For large-scale generalization in Table 2, we generate $100$ new instances.
@@ B.2.3 PARCO Network Hyperparameters
Most hyperparameters are kept similar to Section B.1.3.
@@
Initial Embedding. This layer projects initial raw features to hidden space. For depots, the initial embeddings encode the location $o_{m}$ and the respective vehicle’s capacity $Q_{m}$. For pickup nodes, the initial embeddings encode the location and paired delivery nodes’ location. For delivery nodes, the initial embeddings encode the location and paired pickup nodes’ location. Main Encoder. We employ $l=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
@@
Context Embedding. This layer projects dynamic raw features to hidden space. The context is the embedding for the depot states $o_{m}$, current node states, current length, remaining capacity, and number of visited nodes. These features are then employed to update multiple queries $q_{m},~{}m=1,\dots,{M}$ simultaneously. Main Decoder. Similarly to the encoder, we employ the same hidden dimension and number of attention heads for the Multiple Pointer Mechanism.
@@
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP. Note that unlike the encoder layer, which acts between all $M+N$ tokens, communication layers are lighter because they communicate between ${M}$ agents.
@@
We employ the Priority-based Conflict Handler guided by the model output probability for managing conflicts with priority given to the agent whose probability of selecting the conflicting action is the highest (see Section 4.6).
@@ B.2.4 PARCO Training Hyperparameters
For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions. We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=8$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma & Ba, 2014) with a total batch size $128$ on a single GPU and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $10^{5}$ randomly generated data. Training takes less than 5 hours in our configuration.
@@ B.3.1 Baselines
Gurobi
We implement the mathematical model described above in the exact solver Gurobi (Gurobi Optimization, LLC, 2024) with a time budget of 60 and 600 seconds per instance. However, with both time budgets, Gurobi is only capable of generating solutions to the FFSP20 instances, similar to the findings made by Kwon et al. (2021) for the CPLEX solver.
Random and Shortest Job First (SJF)
The Random and Shortest Job First (SJF) heuristics are simple construction strategies that build valid schedules in an iterative manner. Starting from an empty schedule, the Random construction heuristic iterates through time steps $t=0,\dots T$ and stages $i=1\dots S$ and randomly assigns jobs available at the given time to an idle machine of the respective stage until all jobs are scheduled. Likewise, the SJF proceeds by assigning job-machine pairs with the shortest processing time first. [^2] To obtain results for the heuristics and metaheuristics, we used the implementation of Kwon et al. (2021), provided in the official GitHub repository of the paper: https://github.com/yd-kwon/MatNet
Genetic Algorithm (GA)
Genetic Algorithms are metaheuristics widely used by the OR community to tackle the FFSP (Kahraman et al., 2008). The GA iteratively improves multiple candidate solutions called chromosomes. Each chromosome consists of $S\times N$ real numbers, where $S$ is the number of stages and $N$ is the number of jobs. For each job at each stage, the integer part of the corresponding number indicates the assigned machine index, while the fractional part determines job priority when multiple jobs are available simultaneously. Child chromosomes are created through crossover, inheriting integer and fractional parts independently from two parents. Mutations, applied with a 30% chance, use one of four randomly selected methods: exchange, inverse, insert, or change. The implementation uses 25 chromosomes. One initial chromosome is set to the Shortest Job First (SJF) heuristic solution and the best-performing chromosome is preserved across iterations. Each instance runs for 1,000 iterations.
Particle Swarm Optimization (PSO)
Finally, Particle Swarm Optimization iteratively updates multiple candidate solutions called particles, which are updated by the weighted sum of the inertial value, the local best, and the global best at each iteration (Singh & Mahapatra, 2012). In this implementation, particles use the same representation as GA chromosomes. The algorithm employs 25 particles, with an inertial weight of 0.7 and cognitive and social constants set to 1.5. One initial particle represents the SJF heuristic solution. Like GA, PSO runs for 1,000 iterations per instance.
@@
We benchmark PARCO mainly against MatNet (Kwon et al., 2021), a state-of-the-art NCO architecture for the FFSP. MatNet is an encoder-decoder architecture, which is inspired by the attention model (Kool et al., 2019a). It extends the encoder of the attention model with a dual graph attention layer, a horizontal stack of two transformer blocks, capable of encoding nodes of different types in bipartite graph-like machines and jobs in the FFSP. Kwon et al. (2021) train MatNet using POMO (Kwon et al., 2020).
@@ B.3.2 Datasets
We follow the instance generation scheme outlined in Kwon et al. (2021) sample processing times for job-machine pairs independently from a uniform distribution within the bounds $[2,10]$. For the first three FFSP instance types shown in Table 1 we also use the same instance sizes as Kwon et al. (2021) with $N=20,50\text{ and }100$ jobs and $M=12$ machines which are spread evenly over $S=3$ stages. To test for agent sensitivity in the FFSP, we fix the number of jobs to $N=50$ but alter the number of agents for the last three instance types shown in Table 1. Still, we use $S=3$ for this experiment, but alter the number of machines per stage to $M_{i}=6,8\text{ and }10$, yielding a total of 18, 24 and 30 agents, respectively.
@@
Testing is performed on 100 separate test instances generated randomly according to the above generation scheme.
@@ B.3.3 PARCO Network Hyperparameters
The machines are regarded as the agents in our PARCO framework. As such, their embeddings are used as queries $\bm{q}$ in the Multiple Pointer Mechanism Equation 7, while job embeddings are used as the keys and values. In each decoding step, the machine embeddings are fused with a projection of the time the respective machine becomes idle. Similarly, job embeddings are augmented with a linear transformation of the time they become available in the respective stage before entering the attention head in Equation 6.
@@
We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action is the highest. Formally, priorities $\mathbf{p}_{m}=\log p_{\theta}(a_{m}|x)$ for $m=1,\dots,{M}$.
@@ B.3.4 PARCO Training Hyperparameters
Regarding the training setup, each training instance $i$ is augmented by a factor of 24, and the average makespan over the augmented instances is used as a shared baseline $b^{\text{shared}}_{i}$ for the REINFORCE gradient estimator of Equation 9. We use the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $4\times 10^{-4}$, which we alter during training using a cosine annealing scheme. We train separate models for the environment configurations used in Table 1. We train models corresponding to environments with 20 jobs for 100, with 50 jobs for 150 and with 100 jobs for 200 epochs. In each epoch, we train the models using 1,000 randomly generated instances split into batches of size 50.333Note: to avoid OOMs, For FFSP100 instances, batches are further split into mini-batches of size 25 whose gradients are accumulated.
@@
The following figures visualize the decoding for the machines of a given stage using MatNet and PARCO. As one can see in Figure 5(a), MatNet requires a decoder forward pass for each machine to schedule a job on each of them. In contrast, as detailed in Figure 5(b), PARCO can schedule jobs on all machines simultaneously through its Multiple Pointer Mechanism and Agent Handler, leading to significant efficiency gains.
(a) An FFSP Decoding Step with MatNet
(b) An FFSP Decoding Step with PARCO
Figure 5: Comparison of decoding steps for FFSP with MatNet and PARCO.